OpenStudy (eyust707):

Give a formula for a vector field in the x,y plane with the property that F points in the direction of the origin and has a magnitude inversely proportional to the to the square of the distance to the origin.

OpenStudy (eyust707):



OpenStudy (eyust707):

|F| = (a)(1/x^2+y^2) ???

OpenStudy (eyust707):

both terms under the 1 ^^^

OpenStudy (eyust707):

F must be some f(x,y) i + g(x,y)j

OpenStudy (amistre64):

it is circular yes; a<cos(t),sin(t)> would produce a circle i believe

OpenStudy (amistre64):

lets say: r = a<cos(t),sin(t)> where a is any constant greater than 0 for good luck r' = tangent vector r'' = normal vector

OpenStudy (amistre64):

the normal vector is what i believe you are after, and modify it so it points inward

OpenStudy (eyust707):

i think they just want the equation of a vector field with the property that each vector points towards the origin

OpenStudy (eyust707):

and the magnitude satisfies the requirement

OpenStudy (eyust707):

its only in the x,y plane so how could there be a normal vector?

OpenStudy (amistre64):

a normal vector is just a vector that points out of the curve; which is useful for defining planes

OpenStudy (eyust707):

ohh i thought a normal vector was orthogonal to a plane?

OpenStudy (amistre64):



OpenStudy (amistre64):

orthogonal to the curve, regardless of R^n

OpenStudy (amistre64):

hmm, acceleration here might be inward ... i tend to get those backwards i think

OpenStudy (eyust707):

well lets see if we have a point at say (1,0) and we want the vector to point at the origin we would need F = -i

OpenStudy (eyust707):

nd if we have a point at say (0,1) we would want F= -J

OpenStudy (eyust707):

at (1,1) we would need -xi -yj

OpenStudy (eyust707):

would F = -x i - yj satisfy the origin requirement?

OpenStudy (amistre64):

yes, as long as x and y are functions of t; x(t) and y(t)

OpenStudy (amistre64):

the trick is to define rules for x(t) and y(t) so that they are not so generic

OpenStudy (amistre64):

and x(t) = a cos(t), y(t) = a sin(t) should fit since those define a circle to begin with

OpenStudy (eyust707):

why do we need x as a function of t? it seems as though all we need is an f(x,y) and g(x,y) for this: F = f(x,y)i + g(x,y) j

OpenStudy (amistre64):

because in order to get "the property that F points in the direction of the origin" we have to define a circle, and there are no functions the that define a circle without going to trig ...

OpenStudy (amistre64):

and trig defines a circle in relation to x and y being functions of time

OpenStudy (amistre64):

unless of course im reading it wrong, which is highly probable today :)

OpenStudy (eyust707):

thats the solution from the sol. manual

OpenStudy (amistre64):

ahh, ok. and its using n for the normal ... let me go over it to see if im on the right track at least :)

OpenStudy (eyust707):

ok so at some point in the x,y plane, the distance to the origin in \[\sqrt{x^2 +y^2}\]

OpenStudy (amistre64):

had similar ideas, but i was overthinking it

OpenStudy (eyust707):

and the inverse is (x^2 + y^2)^-1/2

OpenStudy (eyust707):

hmm i dunno i still gont get how they got that final answer

OpenStudy (amistre64):

lets go at it from this manner: r^2 = x^2 + y^2 is the standard eqaution a circle right? such that r = \(\sqrt{x^2+y^2}\)

OpenStudy (eyust707):

yes..

OpenStudy (amistre64):

r^2 = magnitude but |F| is inversely porportional |F| = 1/r^2 |F| = k/(x^2+y^2) is what it reads to me so far

OpenStudy (eyust707):

ohhh! ok

OpenStudy (eyust707):

k being some constant correct?

OpenStudy (amistre64):

yes, a constant of variation if i recall it right

OpenStudy (amistre64):

the normal to function is what they appear to define next, at least the unit normal vector

OpenStudy (eyust707):

so why do they use -x/sqrtx^2+y^2 i -y/sqrtx^2+y^2j to represent the direction instead of just -xi-yj

OpenStudy (amistre64):

i believe it is so that they have a unit vector to work with. I vector that has a magnitude of "1"